OpenStudy (jollypikachu300):

what is the value of the expression 8+4*2 ------ 4*2-4 anwsers are a.12 b.8 c.6 d.4

OpenStudy (mertsj):

Remember: Multiplication before addition

OpenStudy (jollypikachu300):

the line is a fraction

OpenStudy (jollypikachu300):

multiplaction is 4

OpenStudy (mertsj):

Numerator: 8+4*2= 8+8 = 16

OpenStudy (mertsj):

Denominator: 4*2-4= 8-4 = 4

OpenStudy (jollypikachu300):

i will fan and medal

OpenStudy (mertsj):

Can you finish it now?

OpenStudy (jollypikachu300):

yea i can try

OpenStudy (jollypikachu300):

ok so the anwser is 8

OpenStudy (mertsj):

What is the numerator?

OpenStudy (anonymous):

is it 7?